OpenStudy (anonymous):

Convert parametric equation and graph {x=cos^2t,y=sin^2t) Help plzzzzzzzzzzzzzzzzzz

OpenStudy (anonymous):

since \(\cos^2(t)+\sin^2(t)=1\) this is telling you that \(x+y=1\) or \(y=1-x\) a line